# mode=buildnew

### v2

@@ 5.1.1 Benchmark setting
5.2 Results and Discussion
Regarding the dynamics of how the setpoint $b$ decreases: Initially, Equation 27 holds, in which case we do not update the setpoint, even if some but not all the blue curves (corresponding to $R(\cdot)$ components) are below the red curves (components of the setpoint). After all constraints for each component of $R(\cdot)$ are satisfied as described in Equation 26, the setpoint (red curves) also gets adapted to a new value (which we term shrinkage), as a new goal to be reached by the $R(\cdot)$ loss in Equation 10. Note that the shrinkage of setpoint depends on Pareto dominance in Equation 24a, while non-dominance defined an equivalent relation in Definition 2.2. Our hierarchical output feedback optimization training scheme produced the dynamics of $R(\cdot)$ shown in Figure 5 together with the corresponding setpoints $b$ and multipliers. Note that the multiplier $\beta_{y},\gamma_{d},\mu_{recon}$ and their corresponding loss components operate at different numerical ranges, but our controller still manages to drive the different $R(\cdot)$ loss terms down at different scales and rates.
Figure 6 shows the output portrait of $\ell(\cdot)$ versus a component of $R(\cdot)$ of our training scheme in correspondence to Figure 5. With the multi-objective setpoint based model selection criteria in Section 3.2.4, our method admits a multi-objective descent of the selected model compared to the initial output, such that the uncontrolled behavior at the end of iterations can be safely ignored.
From Figure 7, we can observe that the generalization performance of baselines is highly sensitive to hyperparameter selection. In contrast, our method, as supported by the adaptive behaviors visualized in Figure 5, automatically adjusts the multipliers during the training process, leading to a more robust performance with respect to changes of controller hyperparameters and multiplier initial condition $\mu^{(0)}$. Our method does not need hyperparameter searching of the multiplier, thus saves a fair amount of computational resources.
6 Conclusion
This work addresses the issue of combinatorial choice for multidimensional multipliers weighting many terms in neural network loss function. We proposed a probabilistic graphical model for joint multiplier and model parameter optimization with respect to a multi-objective descent of all loss terms. The estimation of model parameter leads to an automatic adjustment scheme for the multipliers, adopting a hierarchical control scheme, which breaks the multi-objective optimization problem into a series of constraint optimization sub-problems. Each sub-problem is configured with a self-adaptive multi-objective setpoint updated via Pareto dominance. A PI-like multiplier controller drives the loss term to satisfy the setpoint constraint for each sub-problem. Our method operates at the timescale of epoch level during training, thus circumvents the need for exhaustive multiplier search and saves tremendous computational resources compared to methods like Bayesian optimization. It also circumvents the excessive memory requirements and heavy computational burden of existing multi-objective deep learning methods. Our method demonstrates robust generalization performance in the PACS domain generalization task with Domain Invariant Variational Autoencoding Ilse et al. (2020), compared to other multiplier choice or scheduling schemes which produce very unstable behavior in the training due to the need for a combinatorial choice of multipliers. Our method is multiplier-free since the multiplier is the output of our controller, and it has robust performance with respect to different multiplier initial values. For future work, we are interested in testing the performance of our hierarchical feedback optimization on other domain generalization methods, as well as on other deep learning fields in the form of Equation 1. Theoretically, it is interesting to validate Conjecture 3.1, derive sufficient conditions, incorporate more mathematical analysis into multidimensional case for $R$, and investigate the potential $\mu$ convergence behavior.
7 Contributions And Acknowledgement
@@
XS proposed the idea, developed the theories and algorithms, implemented the code, designed and carried out the experiments, processed experimental data and generated the figures, wrote the manuscript. NC proposed the PI-like multiplier adaptation in Section 3.2.3, helped XS with PI-like controller implementation, held various discussions, proofread and improved the manuscript. AG initiated the visualization code for Figure 5 and Figure 6, refined by XS, proofread and improved the manuscript. YX developed Corollary 3.2 and corresponding remarks with XS, proofread and improved the constraint optimization part of the theory, discussed on control theory part. CF initiated the benchmark code with refinements from XS, proofread Corollary 3.1. ED and FD tested and improved the code for benchmark, proofread the manuscript. FD further improved the tensorboard visualization of the algorithm. LB proofread Corollary 3.1, tested the benchmark code. DS tested the benchmark code, assisted XS in collecting experimental results. CM supervised the project, offered intensive and in depth proofreads of the manuscript with major quality improvements.

### v3

@@ C.3 Benchmark settings for DIVA
Figure 7: Our automatic multiplier adjustment scheme ensures robust out-of-domain test accuracy when training DIVA with the modified ResNet50 (Gulrajani and Lopez-Paz, 2020) on PACS dataset (testing on domain sketch and training on domain photo, art-painting and cartoon). Each panel (i.e., subplot) corresponds to a training scheme. Each box plot inside a panel corresponds to a specific hyperparameter combination (i.e. multipliers value for baselines and controller hyperparameters for M-HOF-Opt). Inside each box plot, we repeat the experiment with different random seeds corresponding to the dots scatter. In the first panel, we warm up only $\beta_{y}$, $\beta_{d}$ in Equations 45 and 44 to their ultimate value $1.0$ while sampling fixed multiplier $\gamma_{d}$ and $\gamma_{y}$ while fixing $\mu_{recon}=1.0$. In the second panel, additionally, we also warm up $\mu_{recon}$ in Equation 42 to $1.0$ and warm up $\gamma_{d}$ in Equation 46 to the sampled value, and keep $\gamma_{y}$ fixed to be the sampled value. In the third panel, we use fixed constant multipliers without warm-up while keep $\beta=1.0$ and $\mu_{recon}=1.0$ and only sample $\gamma_{d},\gamma_{y}$ combinations. The last two panels correspond to results utilizing our multi-objective hierarchical feedback optimization training method, and we let $\mu$ include all multipliers except $\gamma_{y}=1.0$ in Equation 41. For simplicity, when we write $\mu^{(0)}=1.0$ where $\mu^{(0)}\in\mathbb{R}_{+}^{d}$, we mean each component of $\mu^{(0)}$ equals $1.0$. For reproducibility, see https://github.com/marrlab/DomainLab/tree/mhof.
C.4 Benchmark settings and results for IRM and DIAL
For baseline (named “Fixed Multipliers” in Figure 8), we choose $\mu_{1},\mu_{2}$ in Equation 48 from $0.01,0.1,1.0,10$.
For our method M-HOF-Opt, we choose different controller hyperparameters, with $\mu^{(0)}=1e^{-6},0.001$, $\eta=0.325,0.775$ (uniform between 0 and 1), $\mu^{(clip)}=10,100,1000$.
In Figure 8, each box plot inside the rectangle correspond to a hyperparameter configuration. For baseline “Fixed Multipliers” in Figure 8, this corresponds to $\mu_{1},\mu_{2}$ values in Equation 48. For M-HOF-Opt, each box plot corresponds to different controller hyperparameters.
From Figure 8, we can see that M-HOF-Opt performs robust on out-of-domain generalization against controller hyperparameter changes, while when searching for fixed $\mu_{1},\mu_{2}$ combinations in Equation 48 lead to worse results.
Figure 8: M-HOF-Opt achieves better out-of-domain generalization accuracy compared to using fixed multipliers, when combining regularization loss from Domain Invariant Adversarial Learning (DIAL) and Invariant Risk Minimization (IRM) in Section C.2.2. Each panel (i.e., subplot) corresponds to a training scheme. Each box plot inside a panel corresponds to a specific hyperparameter combination (i.e. multiplier values for baseline “Fixed Multipliers” and controller hyperparameters for M-HOF-Opt). Inside each box plot, we repeat the experiment with different random seeds corresponding to the dots scatter. From the figure, we can see that the baseline performance varies a lot as the multiplier combination changes while M-HOF-Opt remains robust against controller hyperparameter variations. See Sections C.1 and C.4 for benchmark setting.
C.5 Effects of different low level optimizers
In this section, we investigate if our algorithm M-HOF-Opt still works when using different low level (gradient descent) optimization algorithms. Using the same setting as in Section C.4, we obtain Table 1 for changing Adam with AdamW (Loshchilov et al., 2017) and Table 2 for stacking on top of the low level optimization algorithm with CosineAnnealingLR [^3] https://pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.CosineAnnealingLR.html. The results show that M-HOF-Opt is robust against changes of low level optimization algorithms.
Method Mean Std
Warmup 0.767 0.0467
Fixed Multipliers 0.748 0.052
M-HOF-Opt 0.806 0.0456
Table 1: M-HOF-Opt still achieves better out-of-domain generalization accuracy when replacing Adam with AdamW as low level optimizer. For benchmark and experimental setting, see Sections C.1 and C.5.
Method Mean Std
Warmup 0.750 0.0441
Fixed Multipliers 0.7322 0.0570
M-HOF-Opt 0.7873 0.0326
Table 2: M-HOF-Opt still achieves favorable out-of-domain generalization accuracy when stacking CosineAnnealingLR as another hierarchy on top of the low level gradient based optimization algorithm. M-HOF-Opt treats the stacked CosineAnnealingLR and lower level gradient optimization algorithm as the uncontrolled plant. See Sections C.1 and C.5 for benchmark and experimental setting.
Appendix D Computer resources
All experiments were conducted on our internal computation clusters with NVIDIA V100 GPUs.
Appendix E Licenses for Existing Assets
Code
Our implementation leverages several open-source libraries:
In our experiment, we used the Facebook Research version of ResNet from DomainBed (https://github.com/facebookresearch/DomainBed/blob/main/domainbed/networks.py) which is MIT License.
Data
We used the PACS dataset for non-commercial research purposes and we cited the author.
Appendix F Contributions
@@
XS proposed the idea, developed the theories and algorithms, implemented the code, designed and carried out the experiments, processed experimental data and generated the figures, wrote the manuscript. NC proposed the PI-like multiplier adaptation in Section 3.2.5, helped XS with PI-like controller implementation, held various discussions, proofread and improved the manuscript. AG initiated the visualization code for Figure 5 and Figure 6, refined by XS, proofread and improved the manuscript. YX developed Proposition 3.3 and corresponding remarks with XS, proofread and improved the constraint optimization part of the theory, discussed on control theory part. CF initiated the benchmark code with refinements from XS, proofread Proposition 3.2. MW improved the code for DomainLab, especially M-HOF-Opt. ED and FD tested and improved the code for benchmark, proofread the manuscript. FD further improved the tensorboard visualization of the algorithm. LB proofread Proposition 3.2, tested the benchmark code. DS tested the benchmark code, assisted XS in collecting experimental results. CM supervised the project, offered extensive proofreads of the manuscript.